6 years ago Josie was 5 times as old as her
Niece Dianne.How old is Dianne now if her age is one-third of her aunt'#

Please help me.

Answer by ikleyn(52775)

.

Let Dianne's age is x years now.


Then Josie's age is 3x years, according to the condition.


Also, from the condition


    3x-6 = 5*(x-6).        ( <<<---===  6 years ago . . . )


All you need is to solve this equation.


    3x - 6 = 5x - 30

    30 - 6 = 5x - 3x

    24     = 2x

     x     = 24/2 = 12.


ANSWER.  Dianne is 12 years old.

Solved.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.